A 0.770 g mixture containing only sodium chloride and potassium chloride was dissolved in water. It required 102 mL of 0.119 M AgNO3 to completely precipitate all of the chloride present. What is percent composition by mass of sodium chloride and potassium chloride in the mixture?
1
Expert's answer
2020-09-30T04:21:35-0400

Solution.

x = n(NaCl)

y = n(KCl)

M(NaCl) = 58.44 g/mol

M(KCl) = 74.55 g/mol

"n(AgNO3) = C \times V = 0.012 mol"

"58.44 \times x + 74.55 \times y = 0.770"

"x + y = 0.012"

x = 0.0077 mol

y = 0.0043 mol

m(NaCl) = 0.45 g

m(KCl) = 0.32 g

"w(NaCl) = \frac{0.45}{0.77} \times 100 \% = 58.44 \%"

"w(KCl) = \frac{0.32}{0.77} \times 100 \% =41.56 \%"

Answer:

w(NaCl) = 58.44 %

w(KCl) = 41.56 %